Random variable X is distributed Poisson, and P(X 2)-P(X 4). Find P(X 3)
Answer #1

answer is 0.2169

for Poisson distribution with parameter \lambda the probability mass funciont

P(X=x)=exp(-\lambda)\lambda^x/x!

here P(X=2)=P(X=4)

or, exp(-\lambda)\lambda^2/2!=exp(-\lambda)\lambda^4/4!

or,\lambda^2/2=\lambda^4/24

or,\lambda^2=12 and \lambda =sqrt(12)=3.464

now P(X=3)=exp(-3.464)(3.464)^3/3!=0.2169

answer is 0.2169